### v1

# Influencer Identification on Link Predicted Graphs

Abstract
How would admissions look like in a university program for influencers? In the realm of social network analysis, influence maximization and link prediction stand out as pivotal challenges. Influence maximization focuses on identifying a set of key nodes to maximize information dissemination, while link prediction aims to foresee potential connections within the network. Given the complexity of these tasks, especially in large-scale networks, we propose a novel algorithm, The Social Sphere Model, which uniquely combines link prediction techniques with influence maximization strategies to effectively identify future vital nodes in weighted networks. Our approach is tested on two distinct contagion models, offering a promising solution with lower computational demands. This advancement not only enhances our understanding of network dynamics but also opens new avenues for efficient network management and influence strategy development.
link prediction, centrality metrics, vital nodes identification
I Introduction
When a company is launching a new product and wants to use social media for marketing, understanding who to hire for such campaigns becomes of most importance to maximize profit. With the explosion of social media in recent years, a common strategy has been to target certain influential individuals for advertising. These individuals, or “influencers,” become the initial spreaders in different social settings, and through their connections and followers, companies reach greater awareness (and sales) in the entire society. However, social groups are constantly evolving and thus influencers in certain social setting may no longer be the optimal selection in the few months or years until the product is released, which can lower efficiency. It becomes clear, then, that being able to predict future influencers is extremely important.
Within a mathematical setting, one can consider social networks as graphs, where the nodes represent members of the network, and their connections are represented by edges. The study of how nodes can influence a network has been done for many decades, and the reader can refer to recent surveys such as Lü et al.’s listofidentificationtypes in 2016, Pei et al.’s survey2 in 2019, and AbdulAmeer et al.’s survey1 in 2022 to learn about the scope of such work.
In particular, Domingos and Richardson imp1 studied influence through a viral marketing perspective. In 2003, Kempe et al. imp2 formalized this as the influence maximization problem (IMP) of choosing $k$ nodes of greatest influence, proving that a complete solution was NP-hard. However, it is interesting to note that to the best of our knowledge, no mathematical definition of a social Influencer as been proposed in the literature, and thus we shall introduce here what we think can be a useful way of thinking of such important individuals within our society (see Definition 29).
In the present paper we will focus on heuristic measures, revolving around local centralities due to lower time complexity: while global metrics take into consideration more complete information about the graph, they are often computationally expensive. Furthermore, Hu et al. hu claimed that global influence can be approximated by local influence from neighbors of order around three to four, which means local measures may be more cost-effective. As mentioned before, one should keep in mind that social networks are constantly changing. In 2004, Liben-Nowell and Kleinberg linkpredictioncomparison formalized the link prediction problem for social networks as predicting future edges based on knowledge of the current graph. Hasan and Zaki linkpredictionsurvey and Lü and Zhou linkpredictionsurvey2 provided surveys of link prediction results. In our work, we shall focus on local similarity metrics and their quasi-local extensions given by Aziz et al. globalextensions , which are heuristic measures similar to centrality.
We seeks to fill a very important gap in the literature which would allow the understanding of influencer predictions in evolving social networks; that is, the effects of link prediction on heuristic centrality measures. To the best of our knowledge, link prediction similarity metrics have not yet been used with centrality metrics and $k$ node selection algorithms to approximate the effects of time on a dynamic social network.
We shall seek to quantify influence through simple and complex contagion. As stated by Min & Miguel simplecontagion , simple contagion is a model from epidemiology that runs on the assumption that there is a fixed probability of transmission between each pair of vertices. Centola and Macy complexcontagiondefinition defined complex contagion as a model for actions or information with some perceived social cost, considering both personal resistance to influence and forces like peer pressure and social affirmation: a person is infected only if sufficiently many of their neighbors are.
In general terms, easily accepted information spreads as simple contagion whereas more controversial or effort-requiring topics may be complex contagion. Examples of some categorizations can be seen in Table 1 below. We shall begin our study in Section II) by studying link prediction and influencer identification separately on dynamic social networks, and then we shall examine influence through simple and complex contagion models in Section IV.2). By putting the above together we shall present our definition of influencer, detail the specific influence model definitions, and introduce our concrete algorithm, the Social Sphere Model (see Section III).
Information Type Contagion Type Source
Viral Memes
Simple Contagion
Weng et al. memescontagions
Nonviral Memes
Complex Contagion
Weng et al. memescontagions
Spread of Disease
Simple Contagion
Health Behaviors
Complex Contagion
Campbell and Salath’e healthbehaviors
Music
Simple Contagion
Notarmuzi et al. conversationcontroversial
TV Shows
Simple Contagion
Notarmuzi et al. conversationcontroversial
Political/Societal Controversies
Complex Contagion
Notarmuzi et al. conversationcontroversial
Social Movements
Complex Contagion
Centola change
Table 1: Example contagion scenarios and categorizations.
Researchers have noted that the performances of the same set of nodes differ noticeably in the simple and complex contagion (see Figure 1 for an example). For example, Watts and Dodds ltmbehavior found that while the former places a great deal of emphasis on influencers, the latter says away from choosing traditional influencers as initial nodes, especially when considering the relative costs (traditionally influential nodes are likely more expensive hires).
(a) Simple Contagion at time $t=0$
(b) Simple Contagion at time $t=1$
(c) Simple Contagion at time $t=7$
(d) Complex Contagion at time $t=0$
(e) Complex Contagion at time $t=2$
(f) Complex Contagion at time $t=7$
Figure 1: Contagion examples with initial infected node 0 and distances marked on edges (successful infections and successful interactions for complex contagion are marked with red).
In Section V, and through the use of the Social Sphere Model, we observe that influence generally spreads much faster in the simple contagion model than in the complex. We also notice the following:
  • better performance for some parameters in our modified prediction metrics and algorithms when compared to the originals (see Section II);
  • high similarity between our model’s predicted influencers and the ‘true’ future influencers in both selection and collective influence, suggesting the effectiveness of our model; and
  • potential in empirical datasets for identifying influencers overlooked by a static model.
Our paper is organized as follows: in Section II, we explore previous works around heuristic measures in link prediction and influencer identification. We explain the Social Sphere Model and our mathematical definition for influencers in Section III. Our methodology is detailed in Section IV, with evaluation methods in Section IV.1, concrete contagion models in Section IV.2, and experimental setup in Section IV.3. We examine our model’s influence in simple and complex contagion through analysis of our algorithm’s results on both empirical datasets and random graphs in Section V. Finally, we give a summary and future directions for our research in Section VI.
II Social Networks
Social networks are generally modeled using graph theory by expressing the network as a graph $G=(V,E)$, for $V$ the set of vertices (the individuals of the network) and $E$ the set of edges (the connections between individuals). Given such a graph, one may assign a weight function
$w:E\rightarrow[0,1],$
where $w(u,v)$ is the weight of edge $uv$. Such weights may represent a variety of parameters of the social network, and in this paper, we consider weights to be approximations for probabilities of influence transmission, be it information, disease, or more. In particular, we take inspiration from Goyal et al. bernoullidistribution , who took the weights $p(i,j)$ to be the total number of interactions over the total number of trials. In order to study such models, one may assume that the probability of interaction between two nodes during a fixed interval of time is constant, in which case each interval of time can be considered as a Bernoulli trial for temporal datasets.
II.1 Weight functions
Since the geometric distribution counts the number of Bernoulli trials needed to succeed, one can approximate the expected number of intervals needed for an interaction to occur between two nodes $i$ and $j$ as $\frac{1}{w(i,j)}$, the expected value of the geometric distribution with probability $w(i,j)$. These distances are marked on the edges of each graph in Figure 1, and they can be used to approximate the amount of time needed for information to spread along each edge. Formally, we can define them as follows.
Definition 1.
The distance between two vertices $u$ and $v$ though their edge $uv$ is given by the function
$d:E$$\rightarrow \displaystyle\mathbb{R}$
$uv$ $\mapsto$ $\frac{1}{w(u,v)}.$ (1)
For convenience, we will also give some preliminary definitions around neighborhoods, paths, and degree:
Definition 2.
For a graph $G=(V,E)$, define the following:
  • for $v\in V$, the set of all $n$-order neighbors of $v$ is denoted by $N_{n}(v)$;
  • for $u,v\in V$, the set of paths between $u$ and $v$ is denoted by $P(u,v)$;
  • for a path
    $p:(u=v_{0},v_{1},\dots v_{n}=v)\in P(u,v),$
    between $u$ and $v$, we define the values
    $a(p)$ $=$ $\sum_{i=0}^{n-1}d(v_{i},v_{i+1})$ (2)
    $b(p)$ $=$ $\sum_{i=0}^{n-1}w(v_{i},v_{i+1});$ (3)
    $D(u,v)$ $=$ $\min_{p\in P(u,v)}a(u,v).$ (4)
To illustrate the above definitions, consider the graphs in Figure 1. In this setting, one has that
$N_{1}(0)$ $=$ $\{1\}$
$N_{2}(0)$ $=$ $\{3,4\}$
$P(0,2)$ $=$ $\{(0,1,3,2),(0,1,4,2)\}$
$a((0,1,3,2))$ $=$ $2+3+3=8$
$b((0,1,3,2))$ $=$ $\frac{1}{2}+\frac{1}{3}+\frac{1}{3}=\frac{7}{6}$
$D(0,2)$ $=$ 77 77
(achieved by the path $p=(0,1,4,2)$).
Remark 1.
The function $D(u,v)$ calculates the minimum weighted distance between vertices $u$ and $v$.
Definition 3.
Given a graph $G=(V,E)$, let
  • $k_{u}=|N_{1}(u)|$ be the unweighted degree of vertex $u$,
  • $\langle k\rangle=\frac{1}{|V|}\sum_{v\in V}k_{v}$ be the average degree over all vertices in $G$,
  • $s_{u}=\sum_{v\in N(u)}w(u,v)$ be the weighted degree (strength) of vertex $u$, and
  • $\langle s\rangle=\frac{1}{|V|}\sum_{v\in V}s_{v}$ be the average strength over all vertices in $G$.
Coming back to Figure 1’s underlying graph, for instance, we can see that
$k_{0}=1,~{}\langle k\rangle=\frac{1}{5}(1+3+2+2+2)=2,~{}s_{0}=\frac{1}{2}$
and $\langle s\rangle=\frac{1}{5}(\frac{1}{2}+\frac{4}{3}+\frac{2}{3}+\frac{2}{3}+\frac{5}{6})=\frac{4}{5}$.
II.2 Link Prediction
Link prediction is a field that seeks to identify potential future edges in a graph by considering properties of the current graph. There are many types of link prediction strategies, but keeping time complexity in mind we will focus on local methods in this paper, where operating over pairs of nodes will give us similarity scores $s(u,v)$ for all $u\neq v\in V,uv\not\in E$.
The types of link prediction we will be using are path-based, focusing on first-order and second-order neighbors of nodes. We are interested in how a greater penalty for common neighbors with higher degree can affect performance.
We shall give a brief description of the following link prediction models, which shall allow us in particular to see that they can be thought of as variations or extensions of each other:
Common Neighbors. Considering collaboration networks, Newman proposed in commonneighbors that the more common neighbors two nodes share, the more likely they are to form a link in the future, leading to the following definition:
Definition 4.
The common neighbors (CN) similarity score for $u,v\in V$ is
$s_{u,v}^{CN}:=|N(u)\cap N(v)|.$
Murata and Moriyasu weightedlinkprediction amended common neighbors for weighted networks to be dependent on the sum of the weights between the nodes summed over common neighbors leading to the following weighted definition:
Definition 5.
The weighted common neighbors (WCN) similarity score for $u,v\in V$ is
$s_{u,v}^{WCN}:=\sum_{x\in N(u)\cap N(v)}\frac{b(u,x,v)}{2}.$
Local Path (Quasi-Local Common Neighbors). Zhou et al. resourceallocation created the local path metric from the observation that common neighbors as a similarity metric often gives the same scores to many nodes, making the actual rankings ambiguous. They extended the computation of score to include paths of three edges, weighting the longer paths by some $\epsilon$:
Definition 6.
The local path (LP) similarity score for $u,v\in V$ is
$s_{u,v}^{LP}:=|N(u)\cap N(v)|+\epsilon|P_{2}(u,v)|.$
Here, we take $\epsilon=10^{-3}$. Moreover, Aziz et al. globalextensions noted that local path is the quasi-local extension of common neighbors. Extending on the previously mentioned work of Murata and Moriyasu weightedlinkprediction , Bai et al. weightedlocalpath developed a weighted index for local path (weights here are adjusted to match the latter):
Definition 7.
The weighted local path (WLP) similarity score for $u,v\in V$ is
$s_{u,v}^{WLP}:=s_{u,v}^{WCN}+\epsilon\sum_{(u,i,j,v)\in P(u,v)}\frac{b(u,i,j)\cdot b(i,j,v)}{4}.$
Jaccard Coefficient. Liben-Nowell and Kleinberg linkpredictioncomparison derived the Jaccard coefficient from information retrieval, with the following definition:
Definition 8.
The Jaccard coefficient (JC) similarity score for $u,v\in V$ is
$s_{u,v}^{JC}:=\frac{N(u)\cap N(v)}{N(u)\cup N(v)}.$
Within this setting, a penalty is given when the union of the neighborhoods of the nodes is larger. This can be understood as the more friends people have, the less likely that they will both be with a specific common neighbor and then meet.
Resource Allocation. Zhou et al. resourceallocation came up with the resource allocation metric based on the eponymous process where each individual has resources to be split evenly among its neighbors, defined as follows:
Definition 9.
The resource allocation (RA) similarity score for $u,v\in V$ is
$s_{u,v}^{RA}:=\sum_{x\in N(u)\cap N(v)}\frac{1}{|N(x)|}.$
Within this model, a penalty is given to the common neighbors with greater degree on the intuition that a shared neighbor with more neighbors itself is less likely to connect those two specific neighbors when it chooses to introduce people on account of having fewer resources to do so.
Lü and Zhou weightednetworkweakties made a variant of the above resource allocation for weighted graphs, leading to the following:
Definition 10.
The weighted resource allocation (WRA) similarity score for $u,v\in V$ is
$s_{u,v}^{WRA}:=\sum_{x\in N(u)\cap N(v)}\frac{b(u,x,v)}{s(x)}.$
Quasi-Local Resource Allocation. Aziz et al. globalextensions made a quasi-local extension for RA, where one considers paths of three edges in addition to paths of two edges. In other words:
Definition 11.
The quasi-local resource allocation (QRA) similarity score for $u,v\in V$ is
$s_{u,v}^{QRA}:=\left(\sum_{x\in N(u)\cap N(v)}\frac{1}{|N(x)|}\right)+\epsilon\left(\sum_{(u,i,j,v)\in P(u,v)}\frac{1}{|N(i)||N(j)|}\right).$
The weighted quasi-local resource allocation (WQRA) similarity score for $u,v\in V$ is
$s_{u,v}^{WQRA}:=s_{u,v}^{WRA}+\epsilon\sum_{(u,i,j,v)\in P(u,v)}\frac{b(u,i,j)\cdot b(i,j,v)}{s(i)s(j)}.$
A New Resource Allocation Variant: RA22-2- 2. Finally, returning to the intuition that common neighbors with larger degrees contribute less to the similarity between two nodes, we shall introduce here a new similarity metric, novel version of resource allocation. Suppose a node $u$ wants to connect two of its neighbors: there are then $\binom{|N(u)|}{2}$ pairs to choose from, so each pair of its neighbors has a $\frac{1}{\binom{|N(u)|}{2}}$ probability of being linked. For ease of calculation, we approximate this to
$\frac{2}{|N(u)|^{2}}.$
Summing over all common neighbors, we then have the RA22-2- 2 similarity:
Definition 12.
The RA22-2- 2 similarity score for vertices $u,v\in V$ is
$s_{u,v}^{RA2}:=\sum_{x\in N(u)\cap N(v)}\frac{2}{|N(x)|^{2}}.$
The weighted RA22-2- 2 (WRA22-2- 2) similarity score for $u,v\in V$ is
$s_{u,v}^{WRA2}:=\sum_{x\in N(u)\cap N(v)}\frac{b(u,x,v)}{s(x)^{2}}.$
Quasi-Local variant of RA22-2- 2 Inspired by Aziz et al. globalextensions , we also create a quasi-local variant for RA22-2- 2:
Definition 13.
The quasi-local RA22-2- 2 (QR22-2- 2) similarity scorefor $u,v\in V$ is
$s_{u,v}^{QR2}:=\left(\sum_{x\in N(u)\cap N(v)}\frac{2}{|N(x)|^{2}}\right)+\epsilon\left(\sum_{(u,i,j,v)\in P(u,v)}\frac{4}{|N(i)|^{2}\cdot|N(j)|^{2}}\right).$
The weighted quasi-local RA22-2- 2 (WQR22-2- 2) similarity score for $u,v\in V$ is
$s_{u,v}^{WQR2}:=s_{u,v}^{WRA-2}+\epsilon\left(\sum_{(u,i,j,v)\in P(u,v)}\frac{b(u,i,j)\cdot b(i,j,v)}{f(i)^{2}f(j)^{2}}\right).$
In the following sections we shall illustrate the importance of these novel similarity metrics and their use within our Social Sphere Model.
II.3 Node Centrality
Metric Time Complexity (unweighted) Time Complexity (weighted) Type
Degree Centrality
$O(|E|)$degreecomplexity
$O(|E|)$degreecomplexity
local listofinfluencers2
Coreness
$O(|E|)$corenessalgorithm
$O(|E|)$
iterative listofinfluencers2
H-index
$O(|V|\langle k\rangle)$ h-indexcomplexity
$O(|V|\langle k\rangle)$
local listofinfluencers2
LocalRank
$O(|V|\langle k\rangle^{2})$ localrank
N/A
local listofinfluencers2
Clustering Coefficient (and ClusterRank)
$O(|V|\max_{v\in V}k_{v}^{2})$ clusteringcoefftime
$O(|V|\max_{v\in V}k_{v}^{2})$
listofinfluencers2
Closeness
$O(|V||E|)$ timecomplexities
$O(|V||E|+|V|^{2}\log|V|)$ fibheap
path-based listofidentificationtypes , global listofinfluencers2
Betweenness
$O(|V||E|)$ betweennessalgorithm
$O(|V||E|+|V|^{2}\log|V|)$ betweennessalgorithm
path-based listofidentificationtypes , global listofinfluencers2
Eigenvector
$O(|V|^{2}\cdot\text{iteration \#})$ poweriterationmethod
$O(|V|^{2}\cdot\text{iteration \#})$ poweriterationmethod
global, iterative listofinfluencers2
PageRank
$O(|E|\cdot\text{iteration \#})$pagerankcomplexity
$O(|E|\cdot\text{iteration \#})$pagerankcomplexity
iterative listofinfluencers2
Balanced Index
$O(\langle k\rangle V+V\log V)$ ltmmetrics
$O(\langle k\rangle V+V\log V)$
local
Table 2: Time complexities and categorizations of centrality metrics on weighted networks.
Vital node identification is a crucial part of social network analysis, applicable to many field ranging from marketing to epidemiology. Centrality methods, some of the most common of these processes, offer a window into understanding influence dynamics within social networks. Mathematically, centrality methods are based on functions that assign a real-number score to each $v\in V$, essentially ranking the nodes in terms of influence listofidentificationtypes .
The choice of centrality metric often significantly alters the perceptions of who the vital nodes (influencers) are. This section explores common centrality metric definitions, operating principles, and practical implications. We shall give a brief description of the following twelve methods, of which  Table 2 contains a comparative analysis of the time complexities and classifications:
Degree. One of the fastest methods to compute node centrality is by considering its graph theoretical degree, creating a centrality metric from the traditional idea that the more connections a vertex has, the more influential it is.
Definition 14.
The degree centrality of a node $i$ is defined as
$DC_{i}=k_{i}.$
The weighted degree centrality (strength) of a node $i$ is defined as
$S_{i}=\sum_{j\in N(i)}w(i,j)=s_{i}.$
Coreness. In 1983, Seidman kcoreorigin defined the $k$-core as a measure of the network cohesion of a node, where a subgraph is a $k$-core if it has minimum degree at least $k$. Batagelj and Zaveršnik corenessalgorithm define a $k$-core decomposition algorithm in Algorithm 1.
Algorithm 1 $k$-shell decomposition corenessalgorithm
Graph $G$
Sort $V$ by increasing degree
for each $v\in V$ in order do
     core$[v]$ = deg$[v]$
     for $u\in N(v)$ do
         if deg$[u]>$ deg$[v]$ then
              deg$[u]-=1$
              Reorder $u$ in $V$
         end if
     end for
end for
Output list of core values
Later on, Kitsak et al. coreness took the core numbers for each vertex as its corenness value. Eidsaa and Almaas s-core created a weighted version of $k$-core, where node strength is used rather than degree in the decomposition and when a node $u$ is selected (has the minimum strength in the pruned graph), it is in the $\lceil s_{u}\rceil$-core. A similar definition can be taken using $\lfloor s_{u}\rfloor$, and the algorithm is modified accordingly in Algorithm 2 below.
Algorithm 2 weighted $k$-shell decomposition
Graph $G$
Sort $V$ by increasing strength
for each $v\in V$ in order do
     core$[v]=\lfloor s(v)\rfloor$
     for $u\in N(v)$ do
         if deg$[u]>$ deg$[v]$ then
              deg$[u]-=w(u,v)$
              Reorder $u$ in $V$
         end if
     end for
end for
Output list of core values
H-index. Along the same lines of study, but within the academic world, Hirsch h-indexorigin proposed the H-index to measure the impact of a scientist’s research outputs through considering citations per paper. Lü et al. h-index then used it in the context of centrality metrics, introducing the following definition:
Definition 15.
Define the operator $H$ on a finite set $S=\{x_{1},\dots,x_{m}\}$ as the maximum integer $h$ such that there are $h$ elements in $S$ with values $\geq h$. The $n$-order H-index is defined as
$h_{i}^{(n)}=H(h_{j_{1}}^{(n-1)},\dots,h_{j_{k_{i}}}^{(n-1)}),$
where $h_{i}^{(0)}=DC_{i}$.
From the above perspective, the classical H-index can be seen as the $1$-order H-index. Lü et al. h-index proved that the H-index forms a middle ground between a progression from degree centrality to coreness, where the 00-order index is degree and the $\infty$-order index converges to the $k$-core score. Finally, it is interesting to note that Zhao et al. weightedhindex defined the w-lobby index, a metric for weighted networks similar to the H-index, as follows:
Definition 16.
The w-lobby index of a node $u$ is the largest integer $k$ such that |$|\{s(v)\geq k|v\in N(u)\}|\geq k$.
Through the above definitions, we can introduce the $n$-order H-degree, which is defined as
$WH_{i}^{(n)}=WH(wh_{j_{1}}^{(n-1)},\dots,wh_{j_{k_{i}}}^{(n-1)}),$
with $wh_{i}^{(0)}=s(i)$. In this paper, we take the $10$th order H-index.
LocalRank. Aiming to find a more effective local method due to the high time complexities of global and path-based methods, Chen et al. localrank introduced the following extension of degree centrality to more than just first-order neighborhoods:
Definition 17.
The LocalRank score of a node $u$ is
$LR_{u}=\sum_{v\in N(u)}\sum_{w\in N(v)}|N(w)|+|N_{2}(w)|.$
ClusterRank. Watts and Strogatz clusteringcoeff defined the clustering coefficient around the percent of existing links between a node’s neighbors over the total potential amount to quantify how clustered the network is:
Definition 18.
The clustering coefficient of a node $i$ in the directed network $c_{i}$ is defined as
$c_{i}=\frac{|\{(jk\in E|j,k\in N(i)\}|}{k_{i}(k_{i}-1)}.$
Working with the intuition that nodes in more clustered parts of the network may be more redundant and thus have less influence, Chen et al. clusterrank then used the clustering coefficients of nodes to define the following centrality metric:
Definition 19.
The ClusterRank score of a node $i$ is
$f(c_{i})\sum_{j\in N(i)}(k_{j}+1),$
where $f(c_{i})$ is an exponential function of $c_{i}$ of the form $\alpha^{-c_{i}}$ for a fixed constant $\alpha$ (usually $10$).
Remark 2.
This metric considers the clustering coefficient of the neighboring parts of the network: a node whose neighbors are already tightly clustered may be able to spread information within its cluster very quickly, but has more difficulty doing outreach to other communities, whereas nodes that may bridge clusters might be better.
Taking the weighted score to be $f(c_{i})\sum_{j\in N(i)}(s_{j}+1)$ gives a modification for weighted networks.
Closeness. Bavelas introduced in 1948 closenessorigin the idea of closeness centrality as proportional to the reciprocal of the sum of the shortest path lengths from each pair of vertices $u\neq v$. Rephrasing, Freeman closeness defines the concept of closeness centrality as a measure of the relative proximity of a node $p_{k}$ to the rest of the graph $G$, defined as
$C_{C}(p_{k})=\frac{n-1}{\sum_{i=1}^{n}D(p_{k},p_{i})},$
where $V=\{p_{1},\dots,p_{n}\}$ are the vertices of the graph.
In a survey of influencer identification metrics, Lü et al. listofidentificationtypes note that disconnected networks then pose a problem as distances go to infinity, making the nodes indistinguishable by closeness centrality (as they are all then 00). They then give the following modification:
Definition 20.
The closeness centrality of node $i$ is defined as the reciprocal of the harmonic mean of the geodesic distances from $i$ to all other nodes, i.e.
$CC_{i}=\frac{1}{n-1}\sum_{j\neq i}\frac{1}{d_{i,j}}.$
Betweenness. Betweenness centrality considers how critical a node $i$ is to the spread of information, e.g. what percentage of shortest paths going from one node to another must pass through $i$. First formally defined it in 1977 by Freeman betweennessorigin , it can be understood along with an efficient algorithm for computing betweenness centrality, through the work of Brandes betweennessalgorithm :
Definition 21.
The betweennness centrality of a node $i$ is simply
$BC_{i}=\sum_{s,t\neq i,s\neq t}\frac{g_{st}^{i}}{g_{st}},$
where $g_{st}^{i}$ is the number of paths that pass through $i$ among the shortest paths between $s$ and $t$ and $g_{st}$ is the number of shortest paths between $i$ and $j$.
Eigenvector. In eigenvectororigin Landau gave a clarified overview of a previous paper of his detailing the use of eigenvectors to rank chess players, which has been credited with being the first use of eigenvector centrality. Later on, Lü et al. listofidentificationtypes defined the eigenvector centrality scores $EC$ of $G$ as the vector $\vec{x}$ with components $x_{i}=EC_{i}$ such that $\vec{x}:=\frac{1}{\lambda}A\vec{x}$ for the largest eigenvector $\lambda$ of $A$. In other words,
Definition 22.
The eigenvector centrality of a node $i$ is the $i$th component of the eigenvector corresponding to the largest eigenvalue of $A$.
The assumption behind this is that a node’s influence is a linear combination of the influences of its neighbors, weighted by the weights of its outgoing edges.
PageRank. PageRank is a search algorithm presented by Brin and Page pagerankorigin in a paper where they introduced a Google prototype, modeled on the behavior of a bored, link-clicking user. They defined the algorithm as follows:
Definition 23.
The PageRank score $PR(u)$ of each vertex $u$ satisfies the following property:
$PR(u)=(1-d)+d\sum_{v\in N(u)}\frac{PR(v)}{k_{v}},$
where $d$ is a fixed constant in $[0,1]$.
They set $d$ as the damping coefficient, where the $(1-d)$ term represents the probability a user loses interest and goes to a random webpage. For our purposes, we will set $d=1$.
Xing and Ghorbani weightedpagerank modified the PageRank algorithm for weighted graphs:
Definition 24.
The weighted PageRank score $PR(u)$ of each vertex $u$ satisfies the following property:
$WPR(u)=(1-d)+d\sum_{v\in N(u)}\frac{WPR(v)\cdot w(u,v)}{s_{v}},$
where $d$ is a fixed constant in $[0,1]$.
LeaderRank. Inspired by PageRank, Lü et al. leaderrank created the LeaderRank algorithm. Their intention was to include a “ground node” connected to every node in the network, called node $n+1$, and then used the following recursive relationship to solve for their influences:
Definition 25.
At time $t$, the LeaderRank score of node $i$ is defined recursively as $LR_{v}(t)=\sum_{v\in N(u)}\frac{LR_{v}(t-1)}{k_{v}}$.
To differentiate the original nodes from the added one, all nodes from the original graph begin with value $1$ and the leader node begins with value 00. Once the function stabilizes, the leader node’s value gets evenly distributed to the rest of the nodes and deleted. Thus, the LeaderRank scores at the ending time $t_{c}$ are $LR_{u}=LR_{u}(t_{c})+\frac{LR_{n+1}(t_{c})}{|V|}$ for each vertex $u\in V$. Weighted LeaderRank scores can be calculated in the same fashion as weighted PageRank (see Definition 24).
Balanced Index (BI)
Karampourniotis et al. ltmmetrics introduced the Balanced Index metric for complex contagion by taking a weighted combination the metrics of resistance, degree, and neighbors’ degrees. In the context of threshold models, the resistance $r_{v}$ of a node $v$ is defined as a measure of the current threshold of each node, where uninfected nodes $v$ initially begin with $r_{v}=\theta_{v}\cdot s(v)$ and experience decreasing resistance as more of its neighbors become infected. At resistance 00 the node also becomes infected. The neighbors’ degree comes from Karampourniotis et al.’s ltmmetrics indirect drop of resistance metric, which for node $v$ takes the sum over the nodes that $v$ could infect in a single turn (i.e. the nodes $j$ with remaining resistance less than $\frac{w(i,j)}{s(i)}$). The balanced index metric is thus as follows:
Definition 26.
The balanced index score of a node $i$ is
$BI_{i}=a\cdot r_{i}+b\cdot k_{i}+c\sum_{j\in N(i)|r_{j}=1}(k_{j}-1)$
for nonnegative weights $a+b+c=1$.
We can adjust this for weighted graphs by taking
$BI_{i}=a\cdot r_{i}+b\cdot s_{i}+c\sum_{j\in N(i)|r_{j}\leq w(i,j)}(s_{j}-w(i,j)).$
Complex Path Centrality. Seeking to extend the path length measures traditionally used in simple contagion to the wide bridges (reinforcing ties) used in complex contagion paths, Guilbeault and Centola complexpathlength created complex path centrality, based on the following:
Definition 27.
Let $T_{j}$ be the resistance of node $j$. Then,
  • the bridge between nodes $i$ and $j$ (the set of nodes in $N(j)$ that are distance 00 or $1$ away from some node in the neighborhood of $N(i)$) is
    $BW_{i,j}:=\{v|v\in N(j)\land(v\in N(i)\lor\exists u\in N(i):uv\in E)\},$
    with width $W_{i,j}=|B_{i,j}|$ and indicator variable $[W_{i,j}]$ such that $[W_{i,j}]=1$ if $W_{i,j}\geq T_{j}$ and 00 otherwise. A bridge is locally sufficient for contagion spread if $W_{i,j}\geq T_{j}$.
  • $B_{i}$ is the set of nodes $j\neq i$ such that $W_{i,j}>0$,
  • the proportion of locally sufficient bridges is $LB_{i}=\sum_{x\in B_{i}}[W_{i,j}]$,
  • $GEO_{i,j}$ is the shortest path between $i$ and $j$ only through edges between nodes connected by sufficient bridges, and
  • $\phi(GEO_{i,j})$ is the list of vertices in the path $GEO_{i,j}$.
From this, they then defined a complex path from node $i$ to node $j$ as the sequence of sufficiently wide paths from $i$ to $j$. Mathematically, they write it as follows:
Definition 28.
The complex path centrality of a node $i$ is defined as
$CC_{i}=\frac{1}{n-k_{i}}\sum_{i\neq j}|\phi(GEO_{i,j})|.$
II.4 Heuristic Top-$k$ Algorithms
In marketing, one generally finds it prudent to advertise through more than one channel; thus we are concerned with not only finding the single most effective influencer, which centrality metrics aim to do, but with finding the set of $k$ influencers with the greatest collective influence for some $k$ determined by budget.
Transitioning from the single influencer concept to this more practical problem then introduces more complexities. In particular, from Lü et al. listofidentificationtypes , the straightforward approach of taking the highest $k$ nodes by score in a given centrality metric could lead to redundant influences. When information travels relatively unresisted, as in simple contagion (see Table 1), redundancy is often undesirable, as when a single successful interaction suffices to influence an individual, having too much overlap between the chosen nodes’ targets potentially undermines their collective effectiveness and influence. However, in complex contagion, social reinforcement is actually necessary to induce influence cascades, which leads to interesting questions around the straightforward top-$k$ algorithm. We shall thus assess whether the perceived redundancy truly implies reduced performance.
In this section, we explore alternative methodologies to reduce redundancy in the locations of the chosen influencers and modify them to consider centrality metrics as well. We consider the following five algorithms and modify some of them to incorporate centrality metrics, as different centrality metrics may have underlying properties that perform better in tandem with certain algorithms. For the sake of convenience, call these modified algorithms as centrality algorithms.
LIR. To reduce the number of adjacent nodes in the chosen set of $k$, Liu et al. lir_original created the local index rank (LIR) algorithm. They first defined the function $Q(x)$ to be $1$ for $x>0$ and 00 otherwise, then set the $LI$ score of each node $u$ as
$LI(u)=\sum_{v\in N(u)}Q(k_{v}-k_{u}).$
They then focused on the nodes with $LI$ value 00 (disregarding those that simply have no edges) and picked the $k$ nodes from there with largest degree  [^1] In this paper, we modify the original algorithm to incorporate centrality metrics by sorting by the score the node has on a given centrality metric..
Algorithm 3 LIR Algorithm lir_original
$G$, $k$
Find the $LI$ values for all the nodes
Find the list $\{v|LI(v)=0\}$
Sort nodes by degree (decreasing)
Output the top $k$ nodes in the sorted list
LIR-2. Inspired by the LIR method described before, Tao et al. lir2 created an LIR-2 method as a quasi-local extension to second-order neighbors. They define the LIR-2 score of a node $u$ to be
$LI_{2}(u)=\sum_{v\in N(u)}Q(k_{v}-k_{u}),$
as shown in Algorithm 4.
Algorithm 4 LIR-2 Algorithm lir2
$G$, $k$
Find the $LI_{2}$ values for all the nodes
Make sorted list $L$ of sets of nodes with the same $LI_{2}$ values in increasing order
Make an empty list FinalList
for Set $S$ in $L$ do
     Sort nodes in $S$ in decreasing order of degree
     Append to FinalList
end for
Output the top $k$ nodes in FinalList
Joint Nomination. Dong et al. jointnomination proposed the joint nomination algorithm as an algorithm geared towards immunization purposes, where it seeks to select vertices whose removal can help stop influence spread. Their method spans $k$ rounds. In each round, a random node is selected as a nominator, one of its neighbors is randomly selected as conominator, and finally a random common neighbors of the nominator and conominator is selected into the chosen set, as shown in Algorithm 5. For developing our Social Sphere model we shall amend the method of random selection by including probabilities based on centrality metric score and weights  [^2] In this paper, we deviate from the original version of joint nomination by choosing nodes with probability proportional to centrality score, choosing co-nominators with probability proportional to the weight of the edges between them, and choosing a common neighbor (nominee) $i$ with probability $\frac{w(u,i)\cdot w(v,i)}{\sum_{j\in N(u)\cap N(v)}w(u,j)\cdot w(v,j)}$..
Algorithm 5 Joint Nomination jointnomination
$G,k$
Create a set of nominees
Identify a set $N$ of $k$ random nominators
for $u\in N$ do
     Identify set of neighbors $C$
     Randomly pick a co-nominator $v$, where neighbor $i$ is chosen with weight $\frac{w(u,i)}{s(u)}$.
     while $N(u)\cap N(v)$ is $\emptyset$ do
         Remove $v$ from $C$
         Randomly pick another co-nominator $v$
         if $C$ is empty then
              Add a random conominator to the nominees
              Break
         end if
     end while
     if $|N(u)\cap N(v)|>0$ then
         Randomly pick a nominee $n$ from their common neighbors
         Add $n$ to the nominees
     end if
end for
Output nominees
VoteRank. From a different perspective, Zhang et al. voterank proposed VoteRank as an algorithm where influential nodes are chosen by “votes” of the vertices in the graph. There are $k$ rounds of voting. Each node begins with voting power $1$ and score 00 and will divide their voting power among their neighbors, thus changing their scores. In this setting, the node with maximum score is chosen in each round and a resulting penalty given to its neighbors (see Algorithm 6).
Algorithm 6 VoteRank Algorithm voterank
$G$, $k$
Initialize each node $v$ with score $score(v)=0$ and voting ability $va(v)=1$
Create an (empty) list $L$ of chosen nodes
for $i$ in range$(k)$ do
     Reset the scores of all the nodes to 00
     for $u\in V\setminus L$ do
         for $v\in N(u)\setminus L$ do
              Add $1$ to $score(v)$
         end for
     end for
     Pick the node $n$ with greatest score and add it to $L$
     Set the voting power of $n$ to 00
     Decrease the voting power of its neighbors by $\frac{1}{\langle s\rangle}$
end for
Output $L$
VoteRank has been modified in the literature to fit weighted graphs by changing the score to include the importance of having many neighbors (e.g. see Sun et al. wvoterank ). Moreover, VoteRank has also been combined with metrics such as the DIL method (proposed by Liu et al. dil ) and coreness (e.g. see Li et al. dilvoterank and Kumar and Panda kcorevoterank ); inspired by these works, we give a generalization of VoteRank for a given centrality metric $c$ (such as the ones in Section II.3) by setting each node’s voting ability to its score with respect to $c$.
Graph Coloring. Finally, aiming to avoid redundancy, Zhao et al. graphcoloring proposed the graph coloring method as a means of selecting separated vertices. They use the Welsh-Powell algorithm to separate the graph into sets such that each set has a unique color and no nodes in the same set share an edge. For a given centrality metric, they then pick the $k$ nodes with highest score from the largest independent set as their chosen $k$ vertices.
Algorithm 7 Welsh-Powell Algorithm graphcoloring
Graph $G$
Label nodes as $v_{1},\dots,v_{n}$ in descending order according to degree
Let $\pi(v_{1})=1$
for $i$ in $\{1,2,\dots,n-1\}$ do
     Let $C(v_{i+1})=\{\pi(v_{j})|j\leq i,v_{j}\in N(v_{i+1})\}$
     Take the smallest numbered color in the set of colors $C$ not in $C(v_{i+1})$ as $\pi(v_{i+1})$
     $i+=1$
end for
Output the values of the color function $\pi$
Through the extension and modification of the above algorithms, we shall proceed to introduce our Social Sphere Model.
III The Social Sphere Model
Expanding on the previous sections, we shall introduce here our novel model, the Social Sphere Model, an approach designed here to predict future influencers in changing social networks. This model synthesizes key elements of link prediction and top-$k$ influencer algorithms, offering a comprehensive tool for understanding and anticipating changes in social networks. The uniqueness of our approach lies in the integration of these two well-established methodologies to address the dynamic nature of social networks, which, to the best of our knowledge, has not been done before.
III.1 Mathematical Influencers
As mentioned before, one may study social networks through their graph representation, where vertices represent individuals, and their relations are reflected within the edges. In influence maximization, influencers are often defined as those most important to a network, whether structurally or influentially survey2 . In this paper, we shall use a more specific definition of influencer in relation to centrality and influence maximization algorithms.
Definition 29 (Influencer).
A vertex in a network represented by a graph $G$ is an influencer (with respect to some centrality metric) if it is either a $k$-influencer or a single influencer, where
  • a $k$-influencer is a vertex selected through one of the $k$-node selection algorithms with respect to some centrality metric, and
  • a single influencer is the highest scoring vertex with respect to some centrality metric.
III.2 Future Network Prediction
Understanding how to predict the future state of influencers is of upmost importance when considering information spread. Through previously reviewed literature, one is able to do the following:
  • predict future links of our social network, and
  • find the most influential $k$ nodes of a given social network.
We shall combine these and look for future influencers by first predicting the state of the future network. This is achieved by calculating and normalizing link prediction similarity scores into probabilities of transmission, which are then assigned as weights for new edges. Moreover, given the unit of time $t$ we are predicting for, we can find a better predictor for the transmission probability.
Suppose the probability an edge forms between two nodes $u$ and $v$ in one unit of time is $p_{u,v}$. Consider the analogy of a coin flip: an edge forming between two nodes in one unit of time is equivalent to a correspondingly weighted coin landing heads on a single flip. Thus, the probability an interaction has occurred in those $t$ units of time is simply the probability the coin comes up heads some time in its $t$ flips, which is $1-(1-p_{u,v})^{t}$. Similarly, we can take the outcome of the coin flip as whether or not two nodes become connected. Then the probability of connection, and thus transmission of information, is simply $1-(1-p_{u,v})^{t}$, which we take as an updated weight, enabling us to better account for time evolution.
III.3 Future Top-$k$ Nodes Prediction
Our algorithm for future influencer prediction (see Algorithm 8) takes a graph and, given a target $t$ units of time in the future, predicts future connections with a normalized similarity metric. It then applies a top-$k$ algorithm to identify potential future influencers.
Algorithm 8 Future Top-$k$ Nodes Prediction
Graph $G$, Normalized Similarity Metric $M$, Top-$k$ Algorithm $c$, $t$
For vertices $u\neq v$ with $uv\not\in E$, calculate $s^{M}_{u,v}$.
for $u,v$ such that $s^{M}_{u,v}\neq 0$ do
     Add edge $uv$ to $G$ with weight 1($1-(1-s^{M}_{u,v})^{t}$ and distance 11($\frac{1}{1-(1-s^{M}_{u,v})^{t}}$.
end for
Identify the top $k$ nodes on $G$ using $c$.
In what follows we shall first described the methods developed in order to introduce the Social Sphere Model and its main algorithm for future Top-$k$ Nodes Prediction, and then give an analysis of the results one may obtain through the model, as well as describe some of the applications and future directions.
IV Methods
In order to illustrate the different uses of our model and our algorithm, we shall test our algorithm on both a random graph and an empirical dataset: given a graph $G$, we form a training graph from a random subset of existing edges of given size, predict the future state of the training graph, and identify influencers on those predicted graphs. For each graph and parameter tested, we repeated the process ten times, averaging the data. We modelled a dynamic network by taking the training graph as the current configuration and the original graph $G$ as the future network, thus allowing us to assess the efficacy of the influencers found in terms of their ability to “infect” nodes over time when set as the initial spreaders in the graph $G$.
IV.1 Evaluation Methods
Our testing focused on two key aspects:
  • (I)
    the effectiveness of the predicted influencers compared to those found in the actual network, and
  • (II)
    the success of our predictions in identifying true influencers.
To measure the effectiveness mentioned in (I), we quantify influence as the percentage of nodes infected at each unit of time when the chosen influencers are set as initial spreaders on $G$ and an influence model (see Section IV.2) is run. The results are then presented visually in the form of line graphs.
From a different perspective, we describe the predictive success mentioned in (II) as the similarity between our predicted influencers (and their influence) and the influencers found on $G$ (the “true” influencers), quantified in two different ways:
  1. 1.
    accuracy, the fraction of chosen influencers on the predicted graph that are also chosen by the same methods of influencer identification on the original, and
  2. 2.
    mean squared error, a function from statistics that evaluates the differences between the influence of the influencers chosen on the predicted graphs and the original graph; for a given prediction metric and influencer identification algorithm, the mean squared error is
    $\frac{1}{r}\sum_{i=0}^{r}(O(t)-P(t))^{2},$
    where $P(t)$ is the fraction of infected nodes at time $t$ when the influence of the influencers chosen on the predicted graph are evaluated on the original graph and $O(t)$ is defined similarly for the original graph.
IV.2 Contagion Models
To account for different types of information, we evaluate influence according to two simplified models for simple contagion and complex contagion.
Simple Contagion. In general, simple contagion can be understood as an “infection” of information, where an uninfected individual $j$ becomes infected with probability $p$ once news is passed along to them from an infected individual $i$, in a simple percolation model. For simplicity’s sake, we shall take $p=1$ (following SRI models). Then consider our definition of distance as the expected units of time transmission takes; with this model, influence thus can be approximated as transmitting across each edge $uv$ in $d(u,v)$ units of time. Thus, we can see the following:
Definition 30.
In a Simple Contagion model, at time $t$, the infected set $I(t)$ is defined as follows:
$I(t)=\{v\in V|\exists u\in I(0):D(u,v)\leq t\}.$
Complex Contagion. Complex contagion is usually described by threshold models, where each node is given a resistance $\theta_{v}$ and is activated according to a function
$f:I\rightarrow[0,1]$
(see e.g. Shakarian et al. icmltm for further details). While resistance often varies depending on the individual, in this paper we set it to be a fixed percentage of each vertex’s strength, as considering how perceived plurality can lead to adoption, individuals with larger neighborhoods may have more resistance and similarly the converse may happen.
Definition 31.
In a Complex Contagion model, at time $t$, the infected set $I(t)$ is defined as follows:
$I(t)=\{v\in V|\sum_{i\in I(t-1)}w(i,v)\geq\theta\cdot s_{v}\},$
where $\theta\in[0,1]$ is a given threshold fraction.
IV.3 Experimental Setup
We tested our algorithm on an Erdös-Rényi random graph and an existing collaboration network dataset. Erdös and Rényi erdosrenyirgraph defined these random graphs by taking two parameters, $n\in\mathbb{Z}$ and $p\in[0,1]$. The graph is created with $n$ vertices, and each pair of vertices is connected by an edge with probability $p$. We used a graph with $n=500$ and $p=0.05$. The empirical network we used is the arXiv General Relativity and Quantum Cosmology collaboration network from a paper by Leskovec et al. cagrqc , found in the Stanford Large Network Dataset Collection snapnets . The network contains $5242$ nodes and $14496$ edges. In the random graph case, we took $k=5$ and in the arXiv dataset, we took $k=25$. Additionally, we tested the influences of single influencers in each scenario as well.
For each of the datasets, we took two parameters: one when $90\%$ of edges are chosen and $t=1$, and the second one when $70\%$ of the edges are chosen and $t=3$. We are assuming that edges occur at around the same rate each unit of time; thus, the former models influencer prediction in the near future and the latter predicts for more distant networks.
This methodology and experimental setup pave the way for our findings, which offer novel insights into the prediction of influencers in dynamic social networks.
V Results
We shall describe here our results obtained though the methods appearing in the previous section, with insightful findings around the efficacy of our model in accurately predicting future influencers and analysis of the comparative performances of various similarity measures, centrality measures, and algorithms. We shall also break down these results and provide a clearer understanding of our model’s capabilities and limitations below.
V.1 Results on a Random Graph
Thorough applications of our model to Erdö-Rényi random graphs, one can see the following:
  • the average MSE for each of our algorithms is low, indicating a high similarity to the performance of influencers chosen on the “future” graph,
  • the average accuracy differed depending on algorithm and Social Sphere Model parameters, but is often quite high, and
  • the influence of influencers from the predicted graphs is comparable to that of the original and often performes better than influencer identification done on their training graphs.
Prediction Metric Complex Simple Overall
Common Neighbors 0.10224 9.77175E-05 0.05117
Jaccard 0.10427 0.00011 0.05219
Local Path 0.09998 0.00020 0.05009
Quasi-Local RA 0.11917 0.00019 0.05968
Quasi-Local RA22-2- 2 0.10407 0.00020 0.052134
RA22-2- 2 0.10120 0.00010 0.05065
Resource Allocation 0.12174 9.47277E-05 0.06092
Overall 0.10753 0.00014 0.05383
Table 3: Average MSE over all top-$k$ algorithms for link prediction metrics using $70\%$ training graphs from a $500$-node Erdös-Rényi graph.
Prediction Metric Complex Simple Overall
Common Neighbors 0.10127 7.30107E-05 0.05067
Jaccard 0.09809 7.47762E-05 0.04908
Local Path 0.07243 0.00020 0.03632
Quasi-Local RA 0.07870 0.00020 0.03945
Quasi-Local RA-2 0.07062 0.00020 0.03541
RA-2 0.10448 7.62052E-05 0.05228
Resource Allocation 0.10244 7.05446E-05 0.05126
Overall 0.08972 0.00013 0.04492
Table 4: Average MSE over all top-$k$ algorithms for link prediction metrics using $90\%$ training graphs from a $500$-node Erdös-Rényi graph.
Algorithm
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Centrality VoteRank
0.32
0.3467
0.3183
0.3567
0.345
0.345
0.3533
0.3407
Graph Coloring
0.0167 0.0167
0.01
0.01
0.01
0.0167 0.0167
0.01381
Joint Nomination
0.0117
0.0183
0.0117
0.0183
0.0133
0.025
0.0117
0.0157
k Highest
0.4567
0.4617
0.4183
0.4267
0.4417
0.4833
0.465
0.4505
LIR
0.2833
0.3267
0.28
0.3367
0.395
0.3633
0.3
0.3264
LIR-2
0.34
0.36
0.34
0.38
0.44 0.44
0.38
0.3829
Single Influencer
0.375
0.3583
0.3583
0.35
0.3583
0.3667
0.375
0.3631
Random
0
0
0.01 0.01 0.01 0.01 0.01
0.0071
VoteRank
0.3231
0.3477
0.3215
0.3569
0.3446
0.3446
0.3538
0.3418
Overall
0.2529
0.2646
0.2439
0.2634
0.2805
0.2855
0.2667
0.2654
Table 5: Average accuracy over all top-$k$ algorithms for algorithms and link prediction metrics using $70\%$ training graphs from a $500$-node Erdös-Rényi graph.
Algorithm
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Graph Coloring
0.0233
0.0233
0.0600 0.0600 0.0600
0.0233
0.0233
0.0390
Joint Nomination
0.0067
0.0117
0.0100
0.0133
0.0150
0.0133
0.0100
0.0114
k Highest
0.6867
0.6783
0.5867
0.5833
0.5850
0.6867
0.6850
0.6417
LIR
0.5217
0.5367
0.5067
0.5433
0.6033 0.6033
0.5483
0.5519
LIR-2
0.6400
0.6400
0.6400
0.6600 0.6600 0.6600 0.6600
0.6514
Single Influencer
0.6167
0.6000
0.5583
0.5417
0.5417
0.6000
0.5917
0.5786
Random
0.0000
0.0000
0.0100 0.0100 0.0100 0.0100
0.0000
0.0057
VoteRank
0.6
0.62
0.5954
0.6354
0.6308
0.62
0.6277
0.6185
Overall
0.4338
0.4361
0.4149
0.4264
0.4345
0.4497
0.4411
0.4338
Table 6: Average accuracy over all top-$k$ algorithms for algorithms and link prediction metrics using $90\%$ training graphs from a $500$-node Erdös-Rényi graph.
Mean Squared Error (MSE) Comparison. We averaged the MSE for all algorithms and their parameters over each prediction metric in both scenarios tested (see Table 3 and Table 4, where the smallest value in each column appears in bold). In each case, the MSE was quite low, averaging around $0.11$ for complex contagion and $0.00014$ for simple contagion. It is interesting to note that in simple contagion, less variance was observed; this is likely due to the deterministic structure of our simplified influence model. However, even in the more realistic complex contagion scenario we also saw a relatively small mean squared error in the fraction of nodes infected, implying that overall the influence plots for the predicted graphs are quite close to that of the future graph.
Accuracy Comparison. We averaged the accuracy for all algorithms over each centrality metric and prediction metric in both scenarios tested (see Table 5 and Table 6, where the largest value(s) in each row appear in bold). In these tables, we can see the overall predicting ability: the highest accuracies in the $70\%$ scenario are around $0.5$ (implying around half of the chosen vertices are the same on average) and are $0.7$ in the $90\%$ scenario. Therefore one has a high accuracy in the overall selection of influencers, meaning that our predicted graphs do not stray far from the true future graph. In particular, notice that there are a total of $500$ vertices in the graph; when chosen randomly there is an expected overlap of $\frac{5}{500}\cdot 5=0.05$ influencers from linearity of expectation, i.e. an accuracy of $0.0001$.
Interestingly, there are also noticeable differences between the overall accuracies of different algorithms. Overall, the selection of the $k$ highest-scoring centrality metrics performed best in each scenario, with VoteRank, LIR, LIR22-2- 2, and single influencer selection (taking the top scorer with respect to each centrality metric) also performing relatively well. Graph coloring and joint nomination do not perform well with respect to accuracy, however, which is to be expected. For the former, note that with the addition of our various predicted edges, the division of the graph into independent sets was likely changed drastically, which then throws off the accuracy of the nodes chosen in the predicted graphs, and for the latter notice that the algorithm itself relies on a semi-random selection, meaning similar selection of vertices is already quite unlikely. However, since we have weighted the probability distributions when selecting for joint nomination, we do see higher accuracy than expected.
Algorithm Comparisons. We compared both the total overall performance differences by algorithm and their relative performances when also considering a prediction metric.
(a) On complex contagion
(b) On simple contagion
Figure 2: Average performance of each algorithm over all graphs as fraction infected over time using $70\%$ training graphs from a $500$-node Erdös-Rényi graph.
(a) On complex contagion
(b) On simple contagion
Figure 3: Average performance of each algorithm over all graphs as fraction infected over time using $90\%$ training graphs from a $500$-node Erdös-Rényi graph.
We can see from Figure 2 and Figure 3 that LIR, LIR22-2- 2, and VoteRank are the best performers overall. Surprisingly, the $k$ highest algorithm is not too shabby either, surpassing joint nomination and graph coloring in each scenario. Possible explanations for this is first, that the purpose of joint nomination lies with disconnecting the graph rather than creating high influence spread, and second, that since some predicted graphs work better with graph coloring than others due to the aforementioned upsetting of the colorings, its average may have been lowered. As expected, random selection and single influencer selection (labeled as ‘None’ in the figures) perform worst overall.
When considering different prediction metrics, we will look only at data from the complex contagion models (see Figure 4 and Figure 5), as the simple contagion data is very close together for each pair of algorithms and prediction metrics, making it less interesting. In most of the graphs, we can see that compared to ‘None,’ the training graph itself, there are a substantial number of predicted graphs that perform consistently better (RA22-2- 2 and QR22-2- 2, for instance). Moreover, the predicted graphs all display influence approaching or, in 4(c), even surpassing the averages for the original graph, suggesting an improvement in prediction when our Social Sphere Model is used.
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 4: Average performance of each algorithm and prediction metric as fraction infected over time using $70\%$ training graphs from a $500$-node Erdös-Rényi graph.
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 5: Average performance of each algorithm and prediction metric as fraction infected over time using $90\%$ training graphs from a $500$-node Erdös-Rényi graph.
V.2 Results on an Empirical Dataset
The application of our algorithm to the arXiv collaboration network led to the following conclusions:
  • our similarity metric RA22-2- 2 has the lowest average MSE overall,
  • the Social Sphere Model may be able to identify latent influencers who perform even better than those in the original graph, and
  • there always exist parameters for our modified centrality algorithms that perform better than the originals.
Prediction Metric Complex Simple Overall
Common Neighbors 0.02641 0.00557 0.01599
Jaccard 0.02282 0.00493 0.01388
Local Path 0.04713 0.00811 0.02762
Quasi-Local RA 0.04289 0.00753 0.02521
Quasi-Local RA-2 0.04081 0.00738 0.02410
RA-2 0.02254 0.00635 0.01444
Resource Allocation 0.02278 0.00543 0.01411
Overall 0.03220 0.00647 0.01933
Table 7: Average MSE over all top-$k$ algorithms for link prediction metrics using $70\%$ training graphs from the arXiv GrQc dataset cagrqc .
Prediction Metric Complex Simple Overall
Common Neighbors 0.01914 0.01368 0.01641
Jaccard 0.01854 0.01334 0.01594
Local Path 0.02607 0.01365 0.01986
Quasi-Local RA 0.02539 0.01392 0.01966
Quasi-Local RA-2 0.02558 0.01371 0.01965
RA-2 0.01804 0.01372 0.01588
Resource Allocation 0.01846 0.01414 0.01630
Overall 0.02160 0.01374 0.01767
Table 8: Average MSE over top-$k$ algorithms for link prediction metrics using $90\%$ training graphs from cagrqc .
Centrality Metrics Common Neighbors Jaccard Local Path Quasi-local RA-2 Quasi-local RA RA-2 Resource Allocation Overall
Balanced Index
0.0151
0.0106
0.0292
0.0260
0.0247
0.0104
0.0116
0.0182
Betweenness
0.0110
0.0094
0.0224
0.0205
0.0194
0.0087
0.0092
0.0144
Closeness
0.0170
0.0138
0.0295
0.0262
0.0249
0.0133
0.0135
0.0197
ClusterRank
0.0174
0.0151
0.0301
0.0296
0.0287
0.0146
0.0149
0.0215
Complex Path Centrality
0.0475
0.0480
0.0442
0.0434
0.0431
0.0486
0.0484
0.0462
Degree
0.0107
0.0097
0.0229
0.0219
0.0195
0.0086
0.0084
0.0145
Eigenvector
0.0109
0.0095
0.0240
0.0223
0.0212
0.0091
0.0098
0.0152
H-index
0.0156
0.0152
0.0318
0.0282
0.0268
0.0150
0.0152
0.0211
k-core
0.0117
0.0103
0.0222
0.0189
0.0187
0.0109
0.0107
0.0148
LeaderRank
0.0108
0.0083
0.0229
0.0211
0.0196
0.0076
0.0080
0.0140
LocalRank
0.0094
0.0080
0.0231
0.0216
0.0205
0.0077
0.0079
0.0140
PageRank
0.0130
0.0087
0.0266
0.0240
0.0237
0.0099
0.0094
0.0165
Overall
0.0158
0.0139
0.0274
0.0253
0.0242
0.0137
0.0139
0.0192
Table 9: Average MSE for centrality metrics over all algorithms and contagion models on $70\%$ training graphs.
Centrality Metrics Common Neighbors Jaccard Local Path Quasi-local RA-2 Quasi-local RA RA-2 Resource Allocation Overall
Balanced Index
0.0132
0.0132
0.0221
0.0221
0.0218
0.0129
0.0132
0.0169
Betweenness
0.0114
0.0107
0.0199
0.0192
0.0198
0.0102
0.0111
0.0146
Closeness
0.0123
0.0117
0.0201
0.0198
0.0197
0.0111
0.0116
0.0152
ClusterRank
0.0120
0.0120
0.0206
0.0202
0.0204
0.0120
0.0118
0.0156
Complex Path Centrality
0.0732
0.0732
0.0376
0.0374
0.0377
0.0733
0.0734
0.0580
Degree
0.0107
0.0102
0.0172
0.0169
0.0170
0.0099
0.0101
0.0131
Eigenvector
0.0077
0.0074
0.0120
0.0117
0.0116
0.0069
0.0072
0.0092
H-index
0.0125
0.0126
0.0182
0.0179
0.0177
0.0122
0.0124
0.0148
k-core
0.0030
0.0029
0.0048
0.0045
0.0046
0.0031
0.0029
0.0037
LeaderRank
0.0115
0.0108
0.0179
0.0175
0.0181
0.0104
0.0105
0.0138
LocalRank
0.0107
0.0109
0.0198
0.0193
0.0195
0.0103
0.0105
0.0144
PageRank
0.0122
0.0116
0.0204
0.0199
0.0199
0.0113
0.0117
0.0153
Overall
0.0159
0.0156
0.0192
0.0189
0.0190
0.0153
0.0155
0.0170
Table 10: Average MSE for centrality metrics over all algorithms and contagion models on $90\%$ training graphs.
Centrality Metric
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Balanced Index
0.3217
0.4166
0.3091
0.3714
0.3691
0.4086
0.4046
0.3716
Betweenness
0.3811
0.3903
0.3549
0.3537
0.3503
0.3937
0.3943
0.3740
Closeness
0.2897
0.3371
0.2720
0.3120
0.3154
0.3463
0.3411
0.3162
ClusterRank
0.2794
0.2731
0.2217
0.1891
0.1806
0.2760
0.2714
0.2416
Complex Path Centrality
0.2171
0.2160
0.1326
0.1320
0.1343
0.2171
0.2166
0.1808
Degree
0.3634
0.3520
0.3400
0.3211
0.3211
0.3594
0.3594
0.3452
Eigenvector
0.2691
0.2154
0.2577
0.2091
0.2057
0.2240
0.2189
0.2286
H-index
0.1846
0.1509
0.0617
0.0571
0.0560
0.1509
0.1497
0.1158
k-core
0.4680
0.4623
0.4417
0.4394
0.4371
0.4634
0.4611
0.4533
LeaderRank
0.3514
0.3491
0.3297
0.3177
0.3154
0.3560
0.3531
0.3389
LocalRank
0.3000
0.2886
0.2200
0.2057
0.2017
0.2977
0.3000
0.2591
PageRank
0.3766
0.3651
0.3503
0.3291
0.3263
0.3674
0.3674
0.3546
Overall
0.3169
0.3180
0.2743
0.2698
0.2678
0.3217
0.3198
0.2983
Table 11: Average overall accuracies for centrality metrics on $70\%$ training graphs from the arXiv GrQc dataset.
Centrality Metric
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Balanced Index
0.4166
0.4869
0.3989
0.4589
0.4571
0.4857
0.4851
0.4556
Betweenness
0.4743
0.4703
0.4366
0.4343
0.4349
0.4709
0.4697
0.4558
Closeness
0.3874
0.3983
0.3714
0.3857
0.3846
0.4006
0.4000
0.3897
ClusterRank
0.2937
0.2743
0.2080
0.1960
0.1903
0.2743
0.2771
0.2448
Complex Path Centrality
0.1680
0.1640
0.1560
0.1566
0.1571
0.1720
0.1680
0.1631
Degree
0.4606
0.4206
0.4229
0.3869
0.3869
0.4223
0.4194
0.4171
Eigenvector
0.3789
0.3131
0.3600
0.2977
0.3029
0.3217
0.3189
0.3276
H-index
0.1834
0.1743
0.0863
0.0869
0.0834
0.1714
0.1726
0.1369
k-core
0.5274
0.5217
0.4909
0.4829
0.4857
0.5211
0.5211
0.5073
LeaderRank
0.4640
0.4326
0.4274
0.4006
0.4006
0.4360
0.4349
0.4280
LocalRank
0.3280
0.3194
0.2246
0.2200
0.2183
0.3194
0.3194
0.2784
PageRank
0.4851
0.4480
0.4497
0.4143
0.4177
0.4509
0.4486
0.4449
Overall
0.3806
0.3686
0.3360
0.3267
0.3266
0.3705
0.3696
0.3541
Table 12: Average overall accuracies for centrality metrics on $90\%$ training graphs from the arXiv GrQc dataset.
Mean Squared Error (MSE) Comparison. From Table 9 and Table 10, one can see that the MSE of the percentage infected is still small, though substantially larger than before, likely due to the larger variation that comes with more vertices and a larger graph. Notably, our new link prediction metric has the smallest average MSE over all the prediction metrics in both scenarios, with it also being the smallest in $75\%$ of the metrics in Table 10; A similar trend is in Table 11 and Table 12, where our new metric has the highest fraction in the $70\%$ case and the second highest in the $90\%$ scenario. These results support our hypotheses, as there is a significant decrease in MSE in the $90\%$ scenario as opposed to the $70\%$ one, this indicating that our link prediction metrics can identify influencers with very similar influence to that of the original, thus effectively showing prediction.
Accuracy Comparison. When one compares the two scenarios described in Table 11 and Table 12, one can see more variation in the influencers chosen in the second scenario due to less information; in the $90\%$ case, the graph is already very similar to the final configuration and link prediction is more accurate, whereas predictions may be less grounded for the latter case. For more specific data, see Appendix A.
(a) On complex contagion
(b) On simple contagion
Figure 6: Average performance of each algorithm over all graphs as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset.
(a) On complex contagion
(b) On simple contagion
Figure 7: Average performance of each algorithm over all graphs as fraction infected over time using $90\%$ training graphs from the arXiv GrQc dataset.
Influence Comparison. In what follows, we shall consider the relative influences of influencers selected by the following methods:
  1. 1.
    the algorithms,
  2. 2.
    the various prediction metrics when controlling for algorithm, and
  3. 3.
    the various centrality metrics when controlling for algorithm.
The average performances of the various algorithms appears in Figure 6 and Figure 7. Notice that, as in the random graph case, VoteRank performs admirably in both simple and complex contagion in each scenario. Intriguingly, the $k$ highest algorithm also seems to do well, particularly in complex contagion. Unlike in Section V.1, LIR and LIR22-2- 2 exhibit sub-optimal efficiency and have effectively swapped with graph coloring and joint nomination, possibly because of the difference in structure of the empirical dataset and the Erdös-Rényi random graph. For instance, the empirical dataset is not fully connected (thus impossible for the fraction infected to be $1$, as there are too many disconnected pieces).
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 8: Average performance of each algorithm and prediction metric as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset in complex contagion.
(a) Graph Coloring
(b) Joint Nomination
(c) $k$ Highest
(d) LIR
(e) LIR22-2- 2
(f) VoteRank
Figure 9: Average performance of each algorithm and prediction metric as fraction infected over time using $70\%$ training graphs from the arXiv GrQc dataset in simple contagion.
When one controls for algorithm and investigate the average influences of the various prediction metrics (see Figure 8 and Figure 9), we can see that
  • in most cases the respective influence plots of the original graph and the various predicted graphs are very close, and
  • influencers found on the predicted graphs are often able to spread their influence to more of the population.
The first observation entails that the Social Sphere Model is able to produce results comparable with influence identification on the original graph. To elaborate on the second, scenarios such as in 7(a), 7(d), and 8(e) depict the influence plots of predicted graphs as eventually plateauing to values above those of the original graph (and the training graph). This suggests that link predicted graphs may sometimes be able to find latent influencers, those that were influential in the past but overlooked in the current graph.
One should note that Xie et al. dynamicedgedecay have previously explained the drawbacks to having a model that only considers a small window of the graph’s historical data: a powerful influencer who is perhaps temporarily inactive may be overlooked in favor of a currently active, less influential individual. Importantly, our Social Sphere Model thus also presents an incorporation of historical data in finding influencers.
Finally, our analysis of the effectiveness of our modifications to the various algorithms showed that for VoteRank, there exist centrality VoteRank metrics whose averages in accuracy or MSE over all the predicted graphs match or surpass VoteRank’s performance in both the $70\%$ and $90\%$ scenarios, namely closeness, complex path, H-index, $k$-core, and LocalRank in the $70\%$ graph and Balanced Index and LeaderRank in the $90\%$ graph.
When comparing the relative influence of the nodes chosen with VoteRank over the various link prediction metrics, the above is also true (see Figure 10, Figure 11, and Appendix A); in particular, betweenness is shown to be one of the best performing metrics in each scenario (None refers to VoteRank) and closeness and complex path centrality also often do almost as well. This is likely because our contagion models are deterministic and are concerned with the distances between vertices, which the three mentioned metrics, being path-based, are geared to do. Similar observations hold for LIR and LIR22-2- 2.
(a) Original Graph
(b) Common Neighbors Predicted Graph
(c) deg2 RA Predicted Graph
(d) Jaccard Predicted Graph
(e) Local Path Predicted Graph
(f) Quasi-Local deg2 ra Predicted Graph
(g) Quasi-Local RA Predicted Graph
(h) Resource Allocation Predicted Graph
Figure 10: Performance of influencers chosen with Centrality VoteRank (the original VoteRank is labelled as None) on predicted graphs formed from $70\%$ training graphs from the arXiv GrQc dataset in complex contagion.
(a) Original Graph
(b) Common Neighbors Predicted Graph
(c) Jaccard Predicted Graph
(d) Local Path Predicted Graph
(e) Quasi-Local RA Predicted Graph
(f) Quasi-Local RA-2 Predicted Graph
(g) RA-2 Predicted Graph
(h) Resource Allocation Predicted Graph
Figure 11: Performance of influencers chosen with Centrality VoteRank (the original VoteRank is labelled as None) on predicted graphs formed from $70\%$ training graphs from the arXiv GrQc dataset in simple contagion.
VI Conclusion
In this research, we introduced the Social Sphere Model, a novel algorithm which utilizes link prediction and influencer identification to forecast future influencers in weighted networks. Our comprehensive analyses on both a random graph and a real world scenario (through the arXiv collaboration network) have demonstrated the algorithm’s efficacy, particularly evident in the low mean squared error and relatively high accuracy, suggesting the Social Sphere Model’s promising prediction capabilities for detecting future influencers. Moreover, our newly proposed similarity metric, RA-$2$, particularly excelled in both mean squared error and accuracy for the $90\%$ graphs. Moreover, our analysis of different link prediction metrics differing mostly in the degree of their denominators poses interesting questions around the optimal metric when applying the Social Sphere Model for different graphs and parameters.
The applications of our algorithm span a wide array of fields from marketing to epidemiology, offering valuable insights for strategic planning and information control:
  • Viral marketing: taking our algorithm over existing social media data and estimating $t$ from the expected time for production can assist in identifying potential future influencers in the network, giving more time (a valuable commodity) for negotiation, scouting, and strategizing (see Section I). The relatively short time complexities of heuristic models are also sometimes valuable in these scenarios. As an analogy, if one wanted to understand how admissions would look like in a university program for influencers, our algorithm would allow one to evaluate the potential of an influencer or group of influencers in the wider network.
  • Better Network Approximation: The already-documented crimeprediction utility of link prediction in approximating incomplete network data further underscores the significance of our approach, especially considering the typically resource-intensive nature of complete data collection.
  • Historical Consideration: Our results show potential in identifying hidden or overlooked influencers, which presents a dynamic aspect that can enhance performance and integrate more information in the search for influencers.
Since our model is inspired by a deterministic simple contagion model, which restricts the differentiation in algorithm performance, we present here an interesting future research path to explore more realistic contagion models and extend the Social Sphere Model further.
Future studies could also investigate a broader range of link prediction metrics beyond the neighborhood-based and quasi-local metrics employed here, like the clustering coefficient clusteringcoeff . Additionally, the concept of iteratively applying link prediction to evolving graphs, as in futurecommonneighbors , presents an intriguing avenue for extensions to our algorithm and potentially offers a more dynamic understanding of network changes over time. Our use of a variable $t$ in Algorithm 8 seeks to accommodate the increase in probability of interaction over time on the already predicted edges, but it does not account for any future edges that might have formed due to the predicted links’ existence. This paper also chose to include all potential edges, but limiting the additions to the highest-scoring pairs of vertices could increase precision by removing an excess of unlikely edges from posing issues for algorithms like the graph coloring method and the LIR methods.
Finally, it would be of much interest to study the implications of the Social Sphere Model on different network types: in particular, towards identifying the most effective combinations of link prediction and influencer identification metrics for specific network types, such as temporal or directed networks.
Acknowledgments. We extend our gratitude to the MIT PRIMES-USA program for their invaluable support and guidance throughout this research. The research of LPS is partially supported by NSF FRG Award DMS- 2152107, a Simons Fellowship, an NSF CAREER Award DMS 1749013 and a Visiting Fellowship at All Souls College, Oxford.
Affiliations.
  1. (a)
    University of Illinois at Chicago, Chicago, USA.
  2. (b)
    Lakeside High School, Seattle, WA 98107, USA.
Appendix A More Data
Figure 12: Average of accuracies for different algorithms on $70\%$ training graphs for the Erdös-Rényi random graph.
Figure 13: Average accuracies for different algorithms on $90\%$ training graphs for the Erdös-Rényi random graph.
Figure 14: Average accuracies for different algorithms on $70\%$ training graphs from the arXiv GrQc dataset.
Figure 15: Average accuracies for different algorithms on $90\%$ training graphs from the arXiv GrQc dataset.
A complete set of simulation data, csv files, and line plots can be found at https://drive.google.com/drive/folders/1WyuwbMddzajawt4nfh7rLrztY99zO6sm?usp=sharing.